sherry is fives times as old as her son. In five years she will only be three times as old as her son. how old is each now
Answer by addingup(3677)


5x+5= 3(x+5)
5x+5= 3x+15
2x= 10
x= 5 The son is 5. And the mother is 5 times the son:
5*5= 25
In 5 years the son will be:
5+5=10 and the mother three times as old:
10*3= 30 The mother is right now 25 and in 5 years she will be 30. It all makes sense, we have the correct answer.